College of Computer & Information Science Spring 2007 Northeastern University Handout 6 CSG250: Wireless Networks 23 February 2005

Spring 2005 Midterm

Instructions.

• This is a 90-minute open-book open-notes exam.

• The exam is divided into two parts. Part I has 5 short-answer questions, for a total of 15 points. Part II has 5 problems, for a total of 35 points. There is also a bonus subproblem worth 4 points.

• You may want to read all the problems through first and attack them in the order that allows you to make the most progress.

• Your grade on any problem will be determined on the basis of the correctness of the solution and the clarity of the description.

• Show your work, as partial credit will be given. If you need extra space, use the blank facing page.

• Best of luck!

Question Score Maximum
1  3
2  3
3  3
4  3
5  3
6  7
7  7
8  5
9  8
10  8
Bonus  4
Total  50

Name:

Part I (5 × 3 = 15 points)

1. In Pulse Code Modulation (PCM) and Delta Modulation (DM), what are the drawbacks of using linear encoding (equally-spaced quantization levels or equal delta steps)? How can these be addressed?

2. Walsh codes have excellent correlation properties. Yet, they are not used as spreading codes for upstream communication (mobile to base station). Provide a plausible justification for this choice.

3. In the Distributed Coordination Function (DCF) protocol of IEEE 802.11, why does a node wait only SIFS time units (after the last data packet reception) before sending an ACK, while waiting DIFS time units before attempting a data transmission. How critical is this in DCF with an RTS/CTS mechanism? In the Point Coordination Function (PCF)?

4. True or False: Interleaving increases the number of bit errors that can be detected or corrected. Justify your answer and provide the motivation for using interleaving in wireless communication.

5. Consider the handoff procedure in cellular systems that is based on relative signal strength with threshold; that is, a mobile switches from one cell to another if (1) the signal at the current BS is sufficiently weak (less than a predefined threshold) and (2) the other signal is stronger than the two. Discuss the drawbacks of this scheme, when the threshold is too low or too high.

Part II

6. (4 + 3 = 7 points) Antennas and propagation

A microwave transmitter with an output of 0.5 W at 2 GHz is used in a transmission system, where both the transmitting and receiving antennas are parabolas, each 1 m in diameter. Suppose the two antennas are directionally aligned and are 10 kms apart.

(a) What is the effective radiated power of the transmitted signal, in W and dB?

(b) What is the available signal power out of the receiving antenna?

7. (2 + 2 + 3 = 7 points) CDMA codes

Consider a CDMA system in which users A and B have codes (1 1 1 -1 -1 1 -1) and (1 -1 -1 -1 1 1 -1), respectively.

(a) Show the output of the transmission from A at the receiver if A transmits a data bit 1 and if B does not transmit.

(b) Show the output of the transmission from A at the receiver if A transmits a data bit 1 and if B transmits a data bit 1, assuming that the received power from A is same as that from B.

(c) Show the output of the transmission from A at the receiver if A transmits a data bit 1 and if B transmits a data bit 1, assuming that the received power from B is twice as that from A. This can be represented by showing the received signal component from A as consisting of elements of magnitude 1 (+1, −1) and the received signal component from B as consisting of elements of magnitude 2 (+2, −2).

8. (5 points) CRC

Prove that the CRC coding technique can detect any burst errors for which the length of the burst is less than or equal to the length of the frame check sequence (number of redundant bits), as long as the most significant and least significant bits of the CRC code are both 1.